OpenStudy (anonymous):

A university bookstore recently sold a wirebound graph-paper notebook for $2.64, and a college-ruled notebook for $3.61. At the start of spring semester, a combination of 50 of these notebooks were sold for a total of $168.86. How many of each type were sold?

hero (hero):

x + y = 50 2.64x + 3.61y = 168.86 x = graph paper note y = college ruled note

OpenStudy (anonymous):

I am not sure what to do next.

hero (hero):

You solve the system. I'm pretty sure you're working with systems of equations.

hero (hero):

I would recommend using the substitution method

OpenStudy (anonymous):

2.64x + 3.61y = 168.86 sorry trying to look it up cause i am still lost

OpenStudy (anonymous):

I thought you had to have two equations to use the sub method

hero (hero):

We do have two equations don't we?

hero (hero):

I could have sworn I posted two equations initially

OpenStudy (anonymous):

so would it be 3.64x 4.61y =218.86 for the new equation

hero (hero):

No. I don't think you should be working on these kinds of problems unless you are alreadyfamiliar with systems of equations.